In an experiment, the initial concentration of NO2 was 0.00650 mol/L. What is the half-life?


1
Expert's answer
2021-03-18T03:01:54-0400

half life t1/2 = "\frac{[A_o]}{2k}"

[Ao] = initial concentration = 0.00650 mol/L

t1/2 = "\frac{0.00650}{2k}"


t1/2 = "\frac{0.00325}{k}"